Three electric charges of 10.0 micro Coulomb each are at the corners of an equilateral triangle with sides of 30 cm. Determine the potential energy of an 8.0 microcoulomb charge located at the center of the triangle?
1
Expert's answer
2021-08-23T12:39:52-0400
"E_p=3k\frac{qQ}{r^2}\\E_p=3(9\cdot10^{9})\frac{3(10\cdot10^{-6})(8\cdot10^{-6})}{(0.3)^2}=72\ J"